OpenStudy (anonymous):

5x+10/x+2 * 2x/4x-10

OpenStudy (anonymous):

\[\dfrac{ 5x+10 }{ x+2 }\times \dfrac{ 2x }{ 4x-10 }=\dfrac{ 5(x+2) }{ x+2 }\times \dfrac{ 2x }{ 2(2x-5) }\]

OpenStudy (anonymous):

Finish that off.

OpenStudy (anonymous):

LOL it annoying when people don use parenthesis :P

OpenStudy (anonymous):

5(x) / (2x-5)

OpenStudy (anonymous):

You got it. Well Done!

OpenStudy (anonymous):

lol I've been looking at this problem for about an hour telling myself it wasn't right. Thank you